I am trying to figure this out, but I don't understand word problems. Can you help me with this?
A drug manufacturing company wants to manufacture a capsule that contains a spherical pill inside. The diameter of the pill is 4mm and the capsule is cylindrical, with hemispheres on either end. The length of the capsule between the two hemispheres is 10mm. Describe how we could find the exact volume the capsule will hold, excluding the volume of the pill. What is the value that you get from your calculation? Why is it important for us to be able to determine the exact volume of that capsule?

Answer by stanbon(48526)

A drug manufacturing company wants to manufacture a capsule that contains a spherical pill inside. The diameter of the pill is 4mm and the capsule is cylindrical, with hemispheres on either end. The length of the capsule between the two hemispheres is 10mm. Describe how we could find the exact volume the capsule will hold, excluding the volume of the pill. What is the value that you get from your calculation? Why is it important for us to be able to determine the exact volume of that capsule?
---------
Put the two hemispheres together and you have a sphere with
Volume = (4/3)pi(2^3) = (32/3)pi cu.mm.
--------------------------
The body of the pill is a cylinder with
Volume = pi(2^2)*10 = 40pi cu.mm.
--------------------------
Pill volume : (152/3)pi cu mm = 50 2/3 pi cu mm
------------------------------
Medicine dosages are given in mm volume.
------------------------------
Cheers,
Stan H